Please help on this question,


The Hypotenuse of a right triangle is 2.7 units long. The longer leg is 1.3 units longer thant the shorter leg. Find the lengths of the sides of the triangle.
All I am able to figure out is the formula to use is a^2 + b^2 = c^2. Please help. I am lost from here.
Thank you in advance,
Cheryl

Answer by checkley71(8403)

A^2+B^2=C^2
X^2+(1.3X)^2=2.7^2
X^2+1.69X^2=7.29
2.69X^2=7.29
X^2=7.29/2.69
X^2=2.71
X=1.6462 ANSWER FOR THE SHORT LEG.
1.3*1.6462=2.14 ANSWER FOR THE LONG LEG.
PROOF.
1.6462^2+2.14^2=2.7^2
2.71+4.58=7.29
7.29=7.29